# Wei Jiang    Jiayu Yang    Yongqi Zhai    Feng Gao    Ronggang Wang

### v9

@@ 3.2 Overall Architecture
3.2.1 ${}^{++}$
The overall architecture of proposed model is illustrated in Fig. 4. This model is named ${}^{++}$ to distinguish it from MLIC, and ${}^{+}$, which are introduced in our conference version (Jiang et al., 2023b). The architecture of ${}^{++}$, as depicted in Fig. 4, incorporates the analysis transform $g_{a}$, synthesis transform $g_{s}$, hyper analysis $h_{a}$, and hyper synthesis $h_{s}$, which are simplified versions of Cheng’20 (Cheng et al., 2020). To reduce complexity, attention modules are removed.
$N$ $M$ $S$ $K$ Entropy Model
$192$ $320$ $32$ $5$ ${}^{++}$($g_{lc,attn},g_{ch},g_{gc,intra},g_{gc,inter}$)
Table 2: Settings of ${}^{++}$ and ${}^{++}$.
Figure 5: Linear Multi-Reference Entropy Model ${}^{++}$. The figure illustrates the process of decoding a slice $\hat{\boldsymbol{y}}^{i}$.
The hyper-parameters and settings of ${}^{++}$ are presented in Table 2. Same to Minnen et al. (Minnen & Singh, 2020), we adopt mixed quantization, which involves adding uniform noise for entropy estimation and utilizing STE (Theis et al., 2017) to ensure differentiability in the quantization process. Gaussian mean-scale distribution is adopted for entropy estimation. For latent representation $\boldsymbol{y}$, the quantization and estimated rate is formulated as:
$\hat{\boldsymbol{y}}=\mathrm{STE}(\boldsymbol{y}-\boldsymbol{\mu})+\boldsymbol{\mu},\$ (4)
@@ 3.2.1 ${}^{++}$
3.2.2 ${}^{++}$
The proposed linear complexity multi-reference entropy model effectively captures channel-wise, local spatial, and global spatial correlations with linear complexity. The linear complexity entropy model is denoted as ${}^{++}$ to distinguish it from MEM, and ${}^{+}$, which are proposed in our conference version (Jiang et al., 2023b). To capture multi-correlations, the proposed ${}^{++}$ consists of four components: channel-wise context module $g_{ch}$, local spatial context module $g_{lc}$, intra-slice global spatial context module $g_{gc,intra}$, and inter-slice global spatial context module $g_{gc,inter}$. In the channel-wise context module, the latent representation $\hat{\boldsymbol{y}}$ is divided into slices $\{\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\cdots\,\hat{\boldsymbol{y}}^{L}\}$ (Minnen & Singh, 2020) along the channel dimension, $L$ is the number of slices. For the $i$-th slice $\hat{\boldsymbol{y}}^{i}$, the channel-wise context module captures the channel-wise context $\boldsymbol{\Phi}_{ch}^{i}$ from slices $\hat{\boldsymbol{y}}^{<i}$. To capture local spatial correlations, checkerboard pattern (He et al., 2021) is employed, where the latent representation $\hat{\boldsymbol{y}}^{i}$ is divided into anchor part $\hat{\boldsymbol{y}}_{ac}^{i}$ and non-anchor part $\hat{\boldsymbol{y}}_{na}^{i}$. $\hat{\boldsymbol{y}}_{ac}^{i}$ is local-context-free. Local spatial context ${\boldsymbol{\Phi}}_{lc}^{i}$ of $\hat{\boldsymbol{y}}_{na}^{i}$ is captured from $\hat{\boldsymbol{y}}_{ac}^{i}$. We propose Overlapped Window-based Checkerboard Attention $g_{lc,attn}$ for better non-linearity and adaptability to capture local spatial contexts. The global contexts $\boldsymbol{\Phi}_{gc}$ of $i$-th slice are extracted from two dimensions: intra-slice contexts $\boldsymbol{\Phi}_{gc,intra}^{i}$, and inter-slice contexts $\boldsymbol{\Phi}_{gc,inter}^{i}$. We propose Intra-Slice Global Context Module $g_{gc,intra}^{i}$ and Inter-Slice Global Context Module $g_{gc,inter}$ to capture such correlations. Since different slices share the similar global similarity (Jiang et al., 2023b; Guo et al., 2022), the global similarity of $\hat{\boldsymbol{y}}^{i-1}$ is employed to predict the global correlations between $\hat{\boldsymbol{y}}^{i}_{ac}$ and $\hat{\boldsymbol{y}}^{i}_{na}$. The inter-slice global context $\Phi^{i}_{gc,inter}$ are extracted from slices $\hat{\boldsymbol{y}}^{<i}$ via the global similarity of slices $\hat{\boldsymbol{y}}^{<i}$. We introduce these modules in the following sections. The structure of ${}^{++}$ is illustrated in Table 2. We use Equation 1 as our loss function and the estimated rate can be formulated as: $\mathcal{R}=\mathcal{R}_{\hat{\boldsymbol{z}}}+\sum^{L}_{i=0}\left(\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{ac}}+\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{na}}\right)$, where
@@ 3.2.2 ${}^{++}$
[…]
@@ 3.4 Checkerboard Attention-based Local Context Module
One limitation of CNN-based local context modules is their fixed weights, which restricts their ability to capture content-adaptive contexts. We argue that context-adaptation is essential due to the vast diversity of images. In transformers (Vaswani et al., 2017; Dosovitskiy et al., 2021; Liu et al., 2021), the attention weight is generated dynamically according to the input, which inspires us to design a transformer-based content-adaptive local context module. The local receptive field can be envisioned as a window, where local spatial contexts are captured by dividing the feature map into windows. Since each symbol is most relevant to the symbols around it, we propose to make the divided windows overlapped. To achieve this, we propose the novel checkerboard attention context module $g_{lc,attn}$. The process of $i$-th slice is taken as an example. Assuming the resolution of the latent representation $\hat{\boldsymbol{y}}^{i}$ is $H\times W$, the stride is set to $1$ to divide $\hat{\boldsymbol{y}}^{i}$ into $H\times W$ overlapped windows and the window size is $K\times K$. To extract local correlations, the attention map of each window is computed at first. Same as the convolutional checkerboard context module, interactions between $\boldsymbol{y}_{ac}^{i}$ and $\boldsymbol{y}_{na}^{i}$ and interactions in $\boldsymbol{y}_{na}^{i}$ are not allowed. An example of the attention mask is illustrated in Fig. 7. Importantly, this attention mechanism does not alter the resolution of each window. Subsequently, A $K\times K$ convolutional layer is utilized to fuse local context information and and match the size of the local context with that of $\boldsymbol{y}^{i}$ before feeding it to a feed-forward network (FFN) (Vaswani et al., 2017). The overall process is similar to standard transformer (Vaswani et al., 2017). The process is formulated as:
@@
Note that our overlapped window-partition is with linear complexity, since the complexity of each window is $O(K^{4})$. The complexity of $g_{lc,attn}$ is $\Omega(K^{4}SL)$, where $L=HW$, $S$ is the channel number of a slice.
@@ 3.5 Linear Complexity Intra-Slice Global Context Module
During the decoding process, it is challenging to determine the global correlations between the current symbol and other symbols due to the inherent encoding-decoding consistency. This is because the current symbol is unknown during decoding. One potential solution is to embed the global correlations into the bit-stream, but this approach introduces additional bits, thereby increasing the overall bit-rate. Furthermore, in order to obtain precise global similarity, it is necessary to calculate the similarity between the current symbol and all other symbols, which consumes a significant number of bits and is impractical to employ in real-world scenarios. Consequently, representing global similarity with a limited number of bits or without the need for additional bits becomes a non-trivial task.
Figure 7: Mask of Shifted Window-based Checkerboard Attention $g_{lc,attn}$. Blue squares are non-anchor part $\hat{\boldsymbol{y}}_{na}$, white squares are anchor part $\hat{\boldsymbol{y}}_{ac}$. Green square and yellow square are two windows.
Figure 8: Cosine similarity in the spatial domain of different slices of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize slices share similar global correlations.
In latent representation $\hat{\boldsymbol{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 (Cheng et al., 2020), as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\boldsymbol{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\boldsymbol{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\boldsymbol{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@
${\boldsymbol{\Phi}}^{i}_{gc,intra}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$ (14)
@@
where $\hat{\boldsymbol{y}}^{i-1}_{na,q},\hat{\boldsymbol{y}}^{i-1}_{ac,k}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i-1}\right)$, $\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i}_{ac}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects (Kayhan & Gemert, 2020; Islam et al., 2019) of images. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2023a) and is employed to enhance the non-linearity.
@@
One drawback of vanilla approach is its quadratic complexity. In Equation 12, the softmax operation specifies the order of tensor calculation. The complexity of ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ is $O(L^{2})$. The quadratic complexity leads to huge GPU memory consumption, longer encoding and decoding time as illustrated in Fig. 12, which makes it hard to employ the vanilla approach for high-resolution image coding. In Equation 12, if $\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\times\hat{\boldsymbol{y}}^{i}_{ac,v}$ is computed first, the overall complexity becomes $O(L)$, which is linear with the resolution. Equation 12 works because $10<\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)<$ 1. The non-negativity makes it can be treated as a learnable similarity metric. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. To solve the quadratic complexity, it is necessary to introduce a new operator $\rho$ which avoids the necessity to compute ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ first in practice while retaining the non-negativity. Efficient attention operation (Shen et al., 2021) is introduced for non-negativity and linear complexity, which employ the softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na}$ in row and the softmax operation on $\hat{\boldmath{y}}^{i-1}_{ac}$ in column.

### v10

@@ 3.1 Motivation
Figure 4: The overall architecture of MLIC++. $\downarrow$ means down-sampling. $\uparrow$ means up-sampling. / means stride equals $1$. Red line is the dataflow during decoding. ${\boldsymbol{x}}$ is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $Q$ is quantization. $AE$ is arithmetic encoding. $AD$ is arithmetic decoding. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$.
@@ 3.2 Overall Architecture
3.2.1 MLIC++
The overall architecture of proposed model is illustrated in Fig. 4. This model is named MLIC++ to distinguish it from MLIC, and MLIC+, which are introduced in our conference version (Jiang et al., 2023). The architecture of MLIC++, as depicted in Fig. 4, incorporates the analysis transform $g_{a}$, synthesis transform $g_{s}$, hyper analysis $h_{a}$, and hyper synthesis $h_{s}$, which are simplified versions of Cheng’20 (Cheng et al., 2020). To reduce complexity, attention modules are removed.
Figure 5: Linear Multi-Reference Entropy Model MEM++. The figure illustrates the process of decoding a slice $\hat{\boldsymbol{y}}^{i}$.
The hyper-parameters and settings of MLIC++ are presented in Table 2. Same to Minnen et al. (Minnen & Singh, 2020), we adopt mixed quantization, which involves adding uniform noise for entropy estimation and utilizing STE (Theis et al., 2017) to ensure differentiability in the quantization process. Gaussian mean-scale distribution is adopted for entropy estimation. For latent representation $\boldsymbol{y}$, the quantization and estimated rate (Theis et al., 2017; Yang et al., 2023; Ballé et al., 2017, 2018, 2020) is formulated as:
$\hat{\boldsymbol{y}}=\mathrm{STE}(\boldsymbol{y}-\boldsymbol{\mu})+\boldsymbol{\mu},$ (4)
@@ 3.2.1 MLIC++
3.2.2 MEM++
The proposed linear complexity multi-reference entropy model effectively captures channel-wise, local spatial, and global spatial correlations with linear complexity. The linear complexity entropy model is denoted as MEM++ to distinguish it from MEM, and MEM+, which are proposed in our conference version (Jiang et al., 2023). To capture multi-correlations, the proposed MEM++ consists of four components: channel-wise context module $g_{ch}$, local spatial context module $g_{lc}$, intra-slice global spatial context module $g_{gc,intra}$, and inter-slice global spatial context module $g_{gc,inter}$. In the channel-wise context module, the latent representation $\hat{\boldsymbol{y}}$ is divided into slices $\{\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\cdots\,\hat{\boldsymbol{y}}^{L}\}$ (Minnen & Singh, 2020) along the channel dimension, $L$ is the number of slices. For the $i$-th slice $\hat{\boldsymbol{y}}^{i}$, the channel-wise context module captures the channel-wise context $\boldsymbol{\Phi}_{ch}^{i}$ from slices $\hat{\boldsymbol{y}}^{<i}$. To capture local spatial correlations, checkerboard pattern (He et al., 2021) is employed, where the latent representation $\hat{\boldsymbol{y}}^{i}$ is divided into anchor part $\hat{\boldsymbol{y}}_{ac}^{i}$ and non-anchor part $\hat{\boldsymbol{y}}_{na}^{i}$. $\hat{\boldsymbol{y}}_{ac}^{i}$ is local-context-free. Local spatial context ${\boldsymbol{\Phi}}_{lc}^{i}$ of $\hat{\boldsymbol{y}}_{na}^{i}$ is captured from $\hat{\boldsymbol{y}}_{ac}^{i}$. We propose Overlapped Window-based Checkerboard Attention $g_{lc,attn}$ for better non-linearity and adaptability to capture local spatial contexts. The global contexts $\boldsymbol{\Phi}_{gc}$ of $i$-th slice are extracted from two dimensions: intra-slice contexts $\boldsymbol{\Phi}_{gc,intra}^{i}$, and inter-slice contexts $\boldsymbol{\Phi}_{gc,inter}^{i}$. We propose Intra-Slice Global Context Module $g_{gc,intra}^{i}$ and Inter-Slice Global Context Module $g_{gc,inter}$ to capture such correlations. Since different slices share the similar global similarity (Jiang et al., 2023; Guo et al., 2022), the global similarity of $\hat{\boldsymbol{y}}^{i-1}$ is employed to predict the global correlations between $\hat{\boldsymbol{y}}^{i}_{ac}$ and $\hat{\boldsymbol{y}}^{i}_{na}$. The inter-slice global context $\Phi^{i}_{gc,inter}$ are extracted from slices $\hat{\boldsymbol{y}}^{<i}$ via the global similarity of slices $\hat{\boldsymbol{y}}^{<i}$. We introduce these modules in the following sections. The structure of MEM++ is illustrated in Table 2. We use Equation 1 as our loss function and the estimated rate (Theis et al., 2017; Yang et al., 2023; Ballé et al., 2018, 2017, 2020) can be formulated as: $\mathcal{R}=\mathcal{R}_{\hat{\boldsymbol{z}}}+\sum^{L}_{i=0}\left(\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{ac}}+\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{na}}\right)$, where
@@ 3.2.2 MEM++
$\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{ac}}$ $=\mathbb{E}\biggl{[}-\log\int_{-0.5}^{0.5}p\biggl{(}{\boldsymbol{y}}^{i}_{ac}+\boldsymbol{u}|\Phi_{h},$ (7)
$\Phi_{ch}^{i},\Phi_{gc,inter}^{i}\biggr{)}d\boldsymbol{u}\biggr{]},$
$\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{na}}$ $=\mathbb{E}\biggl{[}-\log\int_{-0.5}^{0.5}p\biggl{(}{\boldsymbol{y}}^{i}_{na}+\boldsymbol{u}|\Phi_{h},$ (8)
$\Phi_{lc}^{i},\Phi_{ch}^{i},\Phi_{gc,intra}^{i},\Phi_{gc,inter}^{i}\biggr{)}d\boldsymbol{u}\biggr{]},$
@@ 3.3 Channel-wise Context Module
Figure 7: Visualization of the process and complexity of Shifted Window-based Checkerboard Attention $g_{lc,attn}$. Blue squares are non-anchor part $\hat{\boldsymbol{y}}_{na}$, white squares are anchor part $\hat{\boldsymbol{y}}_{ac}$.
Figure 8: Cosine similarity in the spatial domain of different slices of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize slices share similar global correlations. The similarity maps are nearest-neighbor upsampled for visualization.
@@ 3.4 Checkerboard Attention-based Local Context Module
One limitation of CNN-based local context modules is their fixed weights, which restricts their ability to capture content-adaptive contexts. We argue that context-adaptation is essential due to the vast diversity of images. In transformers (Vaswani et al., 2017; Dosovitskiy et al., 2021; Liu et al., 2021), the attention weight is generated dynamically according to the input, which inspires us to design a transformer-based content-adaptive local context module. The local receptive field can be envisioned as a window, where local spatial contexts are captured by dividing the feature map into windows. Since each symbol is most relevant to the symbols around it, we propose to make the divided windows overlapped. To achieve this, we propose the novel checkerboard attention context module $g_{lc,attn}$. The process of $i$-th slice is taken as an example. Assuming the resolution of the latent representation $\hat{\boldsymbol{y}}^{i}$ is $H\times W$, the stride is set to $1$ to divide $\hat{\boldsymbol{y}}^{i}$ into $H\times W$ overlapped windows and the window size is $K\times K$. To extract local correlations, the attention map of each window is computed at first. Same as the convolutional checkerboard context module, interactions between $\boldsymbol{y}_{ac}^{i}$ and $\boldsymbol{y}_{na}^{i}$ and interactions in $\boldsymbol{y}_{na}^{i}$ are not allowed. An example of the attention mask is illustrated in Fig. 7. Importantly, this attention mechanism does not alter the resolution of each window. Subsequently, a $K\times K$ convolutional layer is utilized to fuse local context information and match the size of the local context with that of $\boldsymbol{y}^{i}$ before feeding it to a feed-forward network (FFN) (Vaswani et al., 2017). The overall process is similar to standard transformer (Vaswani et al., 2017). The process is formulated as:
@@
Note that our overlapped window-partition is with linear complexity, since the complexity of each window is $\Omega(K^{4})$. The complexity of $g_{lc,attn}$ is $\Omega(K^{4}SL)$, where $L=HW$, $S$ is the channel number of a slice.
@@ 3.5 Linear Complexity Intra-Slice Global Context Module
During the decoding process, it is challenging to determine the global correlations between the current symbol and other symbols due to the inherent encoding-decoding consistency. This is because the current symbol is unknown during decoding. One potential solution is to embed the global correlations into the bit-stream, but this approach introduces additional bits, thereby increasing the overall bit-rate. Furthermore, in order to obtain precise global similarity, it is necessary to calculate the similarity between the current symbol and all other symbols, which consumes a significant number of bits and is impractical to employ in real-world scenarios. Consequently, representing global similarity with a limited number of bits or without the need for additional bits becomes a non-trivial task. In latent representation $\hat{\boldsymbol{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 (Cheng et al., 2020), as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\boldsymbol{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\boldsymbol{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\boldsymbol{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@
$\hat{\boldsymbol{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\boldsymbol{y}}^{i}_{attn}),$
${\boldsymbol{\Phi}}^{i}_{gc,intra}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$
@@
where $\hat{\boldsymbol{y}}^{i-1}_{na,q},\hat{\boldsymbol{y}}^{i-1}_{ac,k}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i-1}\right)$, $\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i}_{ac}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects (Kayhan & Gemert, 2020; Islam et al., 2019) of images. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024) and is employed to enhance the non-linearity.
@@
One drawback of vanilla approach is its quadratic complexity. In Equation 12, the softmax operation specifies the order of tensor calculation. The complexity of ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ is $O(L^{2})$. The quadratic complexity leads to huge GPU memory consumption, longer encoding and decoding time as illustrated in Fig. 12, which makes it hard to employ the vanilla approach for high-resolution image coding. In Equation 12, if $\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\times\hat{\boldsymbol{y}}^{i}_{ac,v}$ is computed first, the overall complexity becomes $O(L)$, which is linear with the resolution. Equation 12 works because $10<\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)<$ 1. The non-negativity makes it can be treated as a learnable similarity metric. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)\to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. To solve the quadratic complexity, it is necessary to introduce a new operator which avoids the necessity to compute ${{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{a,k}\right)^{\top}$ first in practice while retaining the non-negativity. Efficient attention operation (Shen et al., 2021) is introduced for non-negativity and linear complexity, which employ the softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na}$ in row and the softmax operation on $\hat{\boldmath{y}}^{i-1}_{ac}$ in column.
@@
The process is illustrated in Equation 13. In Equation 13, $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ is employed as the learnable similarity metric, where $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)<$ 1, $10<\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1, which makes $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\\ \textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1. The non-negativity makes $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ can be employed as a similarity metric. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\\ \to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. In addition, the $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)$ is normalized, which makes each element can be treated probability.
Theorem 3.1.
Same as the standard vanilla attention, each row of the implicit similarity matrix ${softmax}_{2}(\hat{\boldsymbol{y}}_{na,q}^{i}){softmax}_{1}(\hat{\boldsymbol{y}}_{na,k}^{i})^{\top}$ sums up to 1 and represents a normalized attention distribution over all positions.
Proof.
It is evident that each row of the the similarity matrix of the standard attention sums up to $1$.
Let ${softmax}_{2}(\hat{\boldsymbol{y}}_{na,q}^{i})=Q\in\mathbb{R}^{L\times C}$, ${softmax}_{1}(\hat{\boldsymbol{y}}_{na,k}^{i})^{\top}=K\in\mathbb{R}^{C\times L}$, where
$Q=\begin{bmatrix}q_{1,1}&q_{1,2}&\cdots&q_{1,C}\\ q_{2,1}&q_{2,2}&\cdots&q_{2,C}\\ \vdots&\vdots&\ddots&\vdots\\ q_{L,1}&q_{L,2}&\cdots\ &q_{L,C}\\ \end{bmatrix},$ (14)
$K^{\top}=\begin{bmatrix}k_{1,1}&k_{1,2}&\cdots&k_{1,L}\\ k_{2,1}&k_{2,2}&\cdots&k_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ k_{C,1}&k_{C,2}&\cdots\ &k_{C,L}\\ \end{bmatrix}.$ (15)
Due to the characteristics of softmax operation, we can obtain $\sum_{i=1}^{C}q_{j,i}=1$, where $1\leq j\leq L$; $\sum_{i=1}^{L}k_{j,i}=1$, where $1\leq j\leq C$.
$QK^{\top}=\begin{bmatrix}\sum_{i=1}^{C}q_{1,i}k_{i,1}&\cdots&\sum_{i=1}^{C}q_{1,i}k_{i,L}\\ \sum_{i=1}^{C}q_{2,i}k_{i,1}&\cdots&\sum_{i=1}^{C}q_{2,i}k_{i,L}\\ \vdots&\ddots&\vdots\\ \sum_{i=1}^{C}q_{L,i}k_{i,1}&\cdots\ &\sum_{i=1}^{C}q_{L,i}k_{i,L}\\ \end{bmatrix}.$ (16)
The sum of $\ell$-th row is taken as an example.
$\textrm{Sum}_{\ell}$ $=\sum_{i=1}^{C}q_{\ell,i}k_{i,1}+\sum_{i=1}^{C}q_{\ell,i}k_{i,2}+\cdots+\sum_{i=1}^{C}q_{\ell,i}k_{i,L}$ (17)
$=\left(q_{\ell,1}k_{1,1}+q_{\ell,2}k_{2,1}+\cdots+q_{\ell,C}k_{C,1}\right)$
$+\left(q_{\ell,1}k_{1,2}+q_{\ell,2}k_{2,2}+\cdots+q_{\ell,C}k_{C,2}\right)+\cdots$
$+\left(q_{\ell,1}k_{1,L}+q_{\ell,2}k_{2,L}+\cdots+q_{\ell,C}k_{C,L}\right)$
$=q_{\ell,1}\underbrace{\sum_{i=1}^{L}k_{1,i}}_{1}+q_{\ell,2}\underbrace{\sum_{i=1}^{L}k_{2,i}}_{1}+\cdots+q_{\ell,C}\underbrace{\sum_{i=1}^{L}k_{C,i}}_{1}$
$=\sum_{i=1}^{C}q_{\ell,i}=1.$